what is the slope of a line perpendicular to a line with an equation 4x-2y=6
Answer by Alan3354(69443)

what is the slope of a line perpendicular to a line with an equation 4x-2y=6
-----
The slope of 4x - 2y = 6 is 2
To find that, solve for y
4x - 2y = 6
2y + 6 = 4x
2y = 4x - 6
y = 2x - 3
---
The slope, m, is the coefficient of in this format.
m = 2
----
The slope of all lines perpendicular to the given line is the negative inverse.
= -1/2